# Establishing Task Scaling Laws via Compute-Efficient Model Ladders

### v1

@@ 3.3 Chaining the two steps
Finally, we chain the two steps together and predict the task accuracy of a model of size trained on tokens. We first predict the task loss with the fitted function in step 1, and then insert this predicted task loss into the fitted function in step 2 to predict the task accuracy.
@@ 4 Results
We first discuss main prediction results where we chain step 1 and 2 together to make end-to-end predictions. Then we discuss detailed results in step 1 and step 2 individually.
Main results.
@@
Figure 2 shows the prediction results on the 7B-4T and 13B-5T models. On four tasks – MMLU, HellaSwag, PIQA, and Social IQa – we predicted the accuracy of the 7B-4T and 13B-5T models within an absolute error of 2 points. We underestimated ARC-Challenge and ARC-Easy due to the compounding error in the two steps – we overestimated the task loss in step 1, and underestimated the task accuracy in step 2. In §5.1, we analyze the ability of the ladder to predict task performance by considering variation between checkpoints. We find that our results here track with the variation analysis (ARC-Easy and ARC-Challenge display higher variance).
| 7B-4T | 13B-5T | |||||||
| Pred | Actual | Error | %Error | Pred | Actual | Error | %Error | |
| MMLU | 48.4 | 49.0 | 0.6 | 1.3% | 51.3 | 51.6 | 0.3 | 0.7% |
| HellaSwag | 82.5 | 81.3 | 1.2 | 1.4% | 85.3 | 83.2 | 2.1 | 2.5% |
| ARC-Challenge | 51.5 | 61.9 | 10.4 | 16.9% | 52.7 | 63.8 | 11.1 | 17.5% |
| ARC-Easy | 76.6 | 84.6 | 8.0 | 9.4% | 77.2 | 87.2 | 10.0 | 11.4% |
| PIQA | 81.2 | 82.0 | 0.8 | 1.0% | 82.1 | 83.0 | 0.9 | 1.1% |
| CommonsenseQA | 75.7 | 72.6 | 3.1 | 4.2% | 77.6 | 74.1 | 3.5 | 4.7% |
| Social IQa | 58.7 | 59.9 | 1.2 | 2.0% | 59.9 | 61.6 | 1.7 | 2.7% |
| OpenBookQA | 44.2 | 49.4 | 5.2 | 10.6% | 44.9 | 48.6 | 3.7 | 7.8% |
| Average | – | – | 3.8 | 5.9% | – | – | 4.2 | 6.1% |
Step 1 results.
Figure 3 shows the function fitting of step 1 on the ladder models, and the prediction for the target models. The function gives an average relative fitting error ranging from 0.2% to 1.2%, which suggests that Equation 1 is a good functional form to describe task losses. When predicting the task loss for 7B-4T and 13B-5T, we have a relative error within 3% on MMLU, HellaSwag, PIQA, and OpenBookQA. Our method underestimates the loss for CSQA , and overestimates on ARC-ChallengeARC-Easy, and Social IQa. Overestimating the task loss can lead us to eventually underestimate the task accuracy.
Step 2 results.
Figure 4 shows the function fitting of step 2 on the ladder models, and the prediction for the target models with their actual task loss. For all tasks, all data points can be well fitted with a shared sigmoidal function, with the average relative fitting error ranging from 0.4% to 2.6%. When predicting for target models, we get within 3% relative error on MMLU, HellaSwag, PIQA, and CommonsenseQA. We overestimated Social IQa (+4.7%, +2.9%), and underestimated ARC-Challenge (-7.6%, -5.5%), ARC-Easy (-4.5%, -6.0%), and OpenBookQA (-8.2%, -6.0%). The underestimations happen when we cannot accurately extrapolate the slope of the sigmoidal function from data points collected from the small ladder models.
5 Analyses
5.1 Which tasks can the model ladder predict?
@@
Some tasks are inherently more challenging to predict reliably at larger model scales. For example, a test set with an inadequate sample size, low-quality test instances or questions that are too difficult for small models to answer may result in a task that is more challenging for the ladder to predict. We anticipate three sources of prediction failure:
@@
High variance in task loss, resulting in less reliable data points for function fitting.
@@
High variance in task accuracy, leading to a higher spread of prediction targets.
@@
Random-chance task accuracy in the ladder models, due to a task being too difficult to observe any signal from small models.
@@
Using the ladder models that have already been trained, we attempt to predict which tasks will exhibit high errors around the prediction target. We use the intermediate checkpoints of the largest ladder model to measure the noise for task loss and task accuracy. Failure due to random-chance accuracy is discussed when predicting multiple-choice tasks in §B.2.
@@
To quantify the variance of the task loss and accuracy, we compute the standard deviation of the last training checkpoints of the largest ladder model, which we denote as . Similar to our calculation of relative error, we can compare between tasks using relative standard deviation (also called the coefficient of variation), defined as:
A task where ladder models exhibit high variance during training will result in a higher standard deviation across adjacent training checkpoints. This may indicate that one should anticipate a higher error when predicting the final task performance at the target scale. To illustrate , we show the intermediate checkpoints for the largest ladder model on OpenBookQA and MMLU in Figure 5, where we find that captures the apparent noise between adjacent training checkpoints.
@@
We calculate the standard deviation for the task loss and task accuracy using the final 10 checkpoints of the largest ladder model in Table 4, alongside prediction errors for the 7B-4T model. Benchmarks with low SD10, such as MMLU and HellaSwag, also resulted in low prediction errors at the target scale, indicating these tasks are easier for the ladder to predict. We also find that the standard deviation over task loss (Loss SD10) is correlated with step 2 accuracy error (Pearson for 7B-4T and for 13B-5T). In practice, the SD10 can be reported alongside task loss and accuracy predictions to explain which benchmarks may have high error before running the target model.
5.2 How much compute is needed for predicting performance for each task?
In this section, we consider the impact of the scale of the model ladder on the prediction for the target model. In general, the prediction is harder if the difference of scale between the ladder and the target models is larger. We explore this trade-off using the 7B-4T model as the target.
In Figure 6, we order the ladder models by compute-FLOPs, and then start from the smallest compute model and progressively add larger compute models for function fitting. We observe that when using a smaller number of FLOPs, the prediction error is significantly worse. For instance, on MMLU, using our full ladder (3.2% compute of the target model), we get 1.3% error. Reducing the number of ladder FLOPs to 0.1% of the target model compute increases the error to 12% which is an order of magnitude higher than the error with the full ladder. This increase in the prediction error is more observable in certain tasks. Interestingly, we see a slight improvement in errors with less compute in ARC-Challenge and ARC-Easy, potentially due to the variance as seen in §5.1.
Figure 8 similarly shows the impact of including models with longer training regimes. Interestingly, we do not see a significant reduction in the prediction error as we include the ladder models trained to higher Chinchilla multipliers555The target model 7B-4T is trained to approximately 28xC.. There is a slight downward trend (noticeable in MMLU), but this dimension has more flexibility.
6 Ablating Design Choices
We made several design choices in our method, e.g., the two-step approach, using the task loss as an intermediate feature. In this section, we explore some alternatives to these choices. Table 6 provides a summary of the design choices and the corresponding prediction errors. An overarching discussion is presented in §6.5 .
6.1 Task loss prediction: The impact of using compute-FLOPs instead of
In step 1, we used model size and training data size to predict task loss. Here, we test if compute-FLOPs can be used directly for task loss prediction in the over-trained regime with the ladder models. We fit a similar power function
where are parameters to fit.
Using compute-FLOPs fails to distinguish compute-optimal and overtrained models: they can have the same compute-FLOPs but different losses (Hoffmann et al., 2022). In principle, using compute-FLOPs is not expressive enough to generalize to overtrained regimes, and therefore we choose not to use it in our main method.
6.2 Task loss formulation: Incorporating both correct and incorrect answers
Our intermediate feature, task loss, only considers the correct choice of each problem. However, task RC accuracy is determined by the losses of both correct and incorrect choices, which has been cited as a major challenge for predicting downstream performance (Schaeffer et al., 2024):
where is the loss over the ’th answer option of the ’th example and is the label of the correct answer. The loss terms may include some normalization, e.g., by the number of characters in each answer or the unconditional answer probability (see Gu et al., 2024).
We explore taking into account incorrect answers by computing the task cross-entropy, which we define as the cross entropy loss over all answer choices by using the terms as logits. In §C.2, we investigate whether this is a suitable intermediate feature in our task prediction framework. We observe that (1) on some tasks, this intermediate quantity predicts the task accuracy better than using the loss of the correct answer alone, but (2) the two-step predictions are overall worse due to severe errors on MMLU and ARC-Challenge. These tasks incur sizable extrapolation errors of 4-5% in step 1, suggesting that a Chinchilla formulation from Equation 1 may not fully capture the scaling behavior of the task cross-entropy. Furthermore, these errors are amplified in step 2, as the accuracy tends to be far more sensitive to changes in the task cross-entropy than the loss of the correct answer.
6.3 Task loss vs. general language modeling loss
Our method used task loss as the intermediate feature to bridge training scale and task accuracy. Here, we ask if we can instead use a general language modeling loss. Language modeling loss on held-out sets has been shown shown to follow the power law (Hoffmann et al., 2022). Since this loss reflects the capability of LMs, we consider if we can map it to task performance.
We experiment with using the language modeling loss on the C4-en validation set (Raffel et al., 2019) as the intermediate feature. Figure 15 shows the function fitting and predictions. Using C4 loss resulted in higher average prediction error on the task we care most about – MMLU. Prediction error is lower on ARC-Challenge and ARC-Easy. We conclude that using C4 loss can benefit certain tasks where task loss does not work well, and decided to use task loss in our main method because of its superiority on predicting MMLU. We discuss the results in more detail in §C.3.
6.4 The impact of using a two-step prediction instead of combining them into one step
Here we ask, do we need an explicit intermediate feature at all? Can we directly predict task accuracy from the training scale with one function? We explore this possibility by combining Equation 1 and Equation 2 into a single parameterized function:
and we merge parameters and into , so that we reduce to 7 free parameters:
By combining into a single parameterized function, we remove the a priori definition of a specific intermediate feature (i.e., the expression no longer carries a specific meaning), while preserving the representation power of the function. This function can be harder to fit, since it has more free parameters, it is not convex, and we cannot use any data points collected from intermediate checkpoints as we did in step 2. We fit this parameterized function with data points from the final checkpoints of the ladder models, using the same optimization method as in step 1.
Figure 16 shows the function fitting and predictions. The prediction error is worse than using the two-step approach on 4 out of 8 tasks, and the function fitting degenerated on CommonsenseQA. This single-step approach is not as robust as the two-step approach. We discuss the results in more detail in §C.4.
6.5 Discussion of design choices for each task
| Design choice | MMLU | HS | ARC-C | ARC-E | PIQA | CSQA | SIQa | OBQA |
| (N, D) task loss (§3) | 1.3% | 1.4% | 16.9% | 9.4% | 1.0% | 4.2% | 2.0% | 10.6% |
| Task loss over all choices (§6.2) | 18.3% | 7.2% | 21.3% | 5.4% | 3.1% | 2.6% | 0.7% | 7.1% |
| C4 loss (§6.3) | 2.2% | 4.5% | 1.4% | 1.2% | 0.7% | 5.8% | 6.2% | 2.0% |
| Single step (§6.4) | 5.6% | 0.5% | 21.7% | 6.1% | 0.7% | 5.1% | 5.4% | 9.0% |
| Design choice | MMLU | HS | ARC-C | ARC-E | PIQA | CSQA | SIQa | OBQA |
| (N, D) task loss (§3) | 0.7% | 2.5% | 17.5% | 11.4% | 1.1% | 4.7% | 2.7% | 7.8% |
| Task loss over all choices (§6.2) | 20.2% | 10.5% | 19.5% | 6.2% | 2.9% | 2.7% | 1.2% | 0.2% |
| C4 loss (§6.3) | 5.0% | 5.7% | 3.9% | 1.9% | 1.5% | 7.0% | 7.6% | 10.4% |
| Single step (§6.4) | 7.1% | 0.8% | 22.6% | 7.9% | 0.6% | 6.6% | 6.8% | 4.7% |
Table 6 compares prediction errors against alternative design choices. Our two-step method has low error for MMLU, PIQA, HellaSwag, and Social IQa, corresponding to the tasks with the lowest variance around task loss and accuracy, as observed in Table 4.
Tasks with a noisy mapping between their intermediate feature and accuracy, such as the high variance for task loss observed in ARC-Challenge, ARC-Easy and OpenBookQA, may lead to a compounded error when using a multi-step prediction approach. This may explain the higher prediction errors for the design choices in Table 6 which directly or indirectly rely on an intermediate representation of the task loss. For ARC-Challenge and OpenBookQA in particular, their high variance for both the intermediate task loss and accuracy make them challenging for the ladder to predict.
HellaSwag and PIQA, the tasks where we observed the least variance, were the easiest to predict across all design choices. Interestingly, this low variance seems to indicate that the combined single step approach may be capable of predicting performance directly.
Finally, MMLU and HellaSwag have large sample sizes – 5x larger than the other tasks in this work. Given their low sample variance and noise around the prediction target, we expect their errors are most representative of the true difficulty of predicting downstream performance. For this reason, we present the two-step design as our main approach.
Recommendations.
@@ 6.5 Discussion of design choices for each task
Based on our experiments, we present the following guidelines to develop task scaling laws for a new overtrained model / new task: Conduct variance analysis as described in §5.1 for the task(s) using the largest ladder model that can be trained based on available compute (we used 1.3B-10xC). Construct the rest of the ladder as described in §2.1 with a wider range of model sizes compared to (but do train the models beyond 1xC), as discussed in §5.2. For tasks that display high variance, especially for the target metric task accuracy, we generally expect to see lower predictability. High variance in task loss can potentially be mitigated by using alternative intermediate features.
@@ 7 Related Work
Kaplan et al. (2020) and Hoffmann et al. (2022) were among the first to postulate the functional form of language modeling losses as a power function of model parameters and data size. The Chinchilla equation, in particular, has become the founding basis of many subsequent scaling law work (Muennighoff et al., 2023; Gadre et al., 2024; Zhang et al., 2024). This line of work focuses on the language modeling loss on a held-out set with similar distribution as the pretraining data, and much are left to be done on understanding how language models scale on downstream tasks.
@@
Scaling predictions for downstream tasks has been explored in Gadre et al. (2024). Instead of making accuracy predictions directly for individual tasks, they compute the average top-1 error over 17 LLM-foundry (MosaicML, 2024) evaluation tasks as a function of cross entropy loss on C4 (Raffel et al., 2019).
Dubey et al. (2024) uses a two-step prediction to first map the the training compute to the negative log-likelihood of the correct answer for a single task in an evaluation benchmark, and then relate the log-likelihood to the task accuracy. Unlike our work, which uses a fixed ladder of small models, they rely on first finding compute-optimal models.
Chen et al. (2024) also employs a two-stage approach for predicting downstream performance, but uses the pre-training loss instead of a task-specific loss as the intermediate step.
Hu et al. (2023) predicts performance on generative tasks by introducing an intermediate “PassUntil” metric, which is the number of repeated decoding required before landing on a correct answer.
@@ 8 Conclusion and Future Work
In this paper, we develop model ladders and task scaling laws and use them to predict task performance of pretrained language models. With a less than 1% of the pretraining compute, we are able to predict the task performance of 7B-4T and 13B-5T models on individual multiple-choice tasks with good accuracy.
In future work, we hope to reduce the noise in task evaluation metrics by increasing the size of these evaluation sets, which can hopefully yield even lower prediction errors. While we focused on the ranked classification (RC) format of multiple-choice tasks in this work, we hope to make accurate predictions when tasks are written in the multiple-choice (MC) format, which more accurately reflects the capability of larger LMs. MC format accuracy is harder to predict because all our ladder models are too small to exhibit meaningful signal for extrapolation; we show some preliminary results in §B.2 but leave this as future work. Finally, we also hope to validate our method on larger models (e.g., at the 70B scale and beyond).
Limitations
Our task performance prediction method is developed with and validated on the OLMo 2 family of models. In principle, the method should be transferrable to any transformer-based language models, though it has not been validated in practice. We focused on predicting performance on multiple-choice tasks written in ranked classification (RC) format, leaving multiple-choice format (MC) and generative tasks as future work.

### v2

@@ 3.3 Chaining the two steps
We chain the two steps by first predicting the intermediate loss with the fitted function in step 1, and then inserting it into the fitted function in step 2 to predict the task accuracy.
@@
Using task loss as the intermediate feature, we get an average absolute error of 3.8 points for 7B-4T, and 4.2 points for 13B-5T. In §C.5, we also show the results of predicting the task accuracy directly from (N, D) in a single step, and observe higher average prediction errors for both target models ( 30 points).
@@ 4 Results
| 7B-4T | 13B-5T | |||||||||||||||
| Design choice | MMLU | HS | ARC-C | ARC-E | PIQA | CSQA | SIQa | OBQA | MMLU | HS | ARC-C | ARC-E | PIQA | CSQA | SIQa | OBQA |
| (N, D) task loss (§3) | 1.3% | 0.3% | 7.0% | 13.3% | 2.0% | 11.7% | 4.3% | 1.3% | 0.2% | 1.2% | 9.4% | 16.0% | 2.7% | 18.5% | 3.6% | 0.9% |
| FLOPs task loss (§C.1) | 4.3% | 2.1% | 2.4% | 5.3% | 2.0% | 12.5% | 5.7% | 0.5% | 5.2% | 2.6% | 3.5% | 7.0% | 2.6% | 18.8% | 5.7% | 1.4% |
| 7B-4T | 13B-5T | |||||||||||||||
| Design choice | MMLU | HS | ARC-C | ARC-E | PIQA | CSQA | SIQa | OBQA | MMLU | HS | ARC-C | ARC-E | PIQA | CSQA | SIQa | OBQA |
| (N, D) task loss (§3) | 1.3% | 1.4% | 16.9% | 9.4% | 1.0% | 4.2% | 2.0% | 10.6% | 0.7% | 2.5% | 17.5% | 11.4% | 1.1% | 4.7% | 2.7% | 7.8% |
| TaskCE (§C.2) | 18.3% | 7.2% | 21.3% | 5.4% | 3.1% | 2.6% | 0.7% | 7.1% | 20.2% | 10.5% | 19.5% | 6.2% | 2.9% | 2.7% | 1.2% | 0.2% |
| C4 loss (§C.3) | 2.2% | 4.5% | 1.4% | 1.2% | 0.7% | 5.8% | 6.2% | 2.0% | 5.0% | 5.7% | 3.9% | 1.9% | 1.5% | 7.0% | 7.6% | 10.4% |
| Single step (§C.5) | 5.6% | 0.5% | 21.7% | 6.1% | 0.7% | 5.1% | 5.4% | 9.0% | 7.1% | 0.8% | 22.6% | 7.9% | 0.6% | 6.6% | 6.8% | 4.7% |
Table 3 compares prediction errors for all design choices.
Task loss.
@@
Figure 2 shows the chained prediction results on the target models using task loss as the intermediate feature. On four tasks – MMLU, HellaSwag, PIQA, and Social IQa – we predict the accuracy within an absolute error of 2 points. For ARC-C and ARC-E, we overestimate the task loss in step 1, and underestimate the task accuracy in step 2. In §5, we analyze the ability of the ladder to predict task performance by considering variation between checkpoints. We find that our results here track with the variation analysis (ARC-E and ARC-C display higher variance).
TaskCE.
LM loss.
Interestingly, the loss on C4-en validation set is a good predictor of task accuracy for several tasks. One possible explanation for this is potential domain overlap between the task and the validation set. We discuss pros and cons of using LM loss in §6. Full results using LM Loss are show in §C.3.
Figure 16 compares absolute and relative errors for all three intermediate features.
5 Analysis: Task predictability with the ladder
@@
Some tasks are inherently more challenging to predict reliably with the ladder models; for example, a test set with an inadequate sample size, low-quality test instances or questions that are too difficult for small models. We anticipate three sources of prediction failure:
@@
High variance in intermediate loss; less reliable data points for function fitting.
@@
High variance in task accuracy; higher spread of prediction targets.
@@
Random-chance task accuracy in smaller ladder models due to task difficulty.
@@
We try to determine which tasks will exhibit high prediction errors, using task loss as the intermediate feature. We use the intermediate checkpoints of the largest ladder model (1B-1xC) to measure the noise for task loss and task accuracy. Failure due to random-chance accuracy is discussed when predicting multiple-choice tasks in §B.2.
@@
We compute the standard deviation of the last training checkpoints of 1B-10xC (). We also compute relative SD (also called the coefficient of variation) to compare between tasks:
A task where ladder models exhibit a higher standard deviation across adjacent evaluated checkpoints potentially indicates a higher prediction error for the target models. To illustrate , we show the intermediate checkpoints for the largest ladder model on OpenBookQA and MMLU in Figure 5, where we find that captures the apparent noise between adjacent training checkpoints.
@@
We present for the 1B-10xC model Table 4, alongside prediction errors for the 7B-4T model. Benchmarks with low SD10 (e.g. MMLU and HellaSwag), also have low prediction errors for the target, indicating these tasks are easier for the ladder to predict. We also find that Loss SD10 is correlated with step 2 accuracy error (Pearson for 7B-4T and for 13B-5T). Thus, SD10 can be reported alongside predictions to explain which benchmarks may have high error before running the target model.
6 Discussion and Recommendations
In §4, we observed that different design choices work well for different tasks. Here, we discuss factors that affect final predictions, and present some guidelines to develop task scaling laws for a new overtrained model or new task (Figure 16 illustrates these findings):
Target variance.
@@ 6 Discussion and Recommendations
Variance analysis is useful for determining if it is feasible to predict the target metrics (both intermediate loss and accuracy). For tasks with high variance, especially for task accuracy, we generally expect to see lower predictability. High variance in a particular intermediate loss can potentially be mitigated with alternative intermediate features. For new tasks, variance analysis should be conducted on the largest ladder model that can be trained with available compute (we used 1B-10xC).
Choice of intermediate features.
It is important for the intermediate feature to be both 1) predictable, and 2) a good predictor for task accuracy. Tasks with a noisy mapping between their intermediate feature and accuracy (e.g., ARC-C, ARC-E and OpenBookQA) may have compounded errors in the two-step prediction. For ARC-C and OpenBookQA in particular, their high variance for both the intermediate task loss and accuracy make them challenging to predict with the ladder. HellaSwag and PIQA, the tasks where we observe the least variance, are the easiest to predict across all design choices, including the combined single step approach. We also note that while LM loss on C4 is a reasonable predictor of task accuracy for several tasks in our setup, this may not hold for other novel downstream tasks on different domains. As a general rule, using a task-specific loss may work across a wider range of downstream tasks.
Number of task instances.
Tasks with fewer instances show inconsistent results. E.g. OpenBookQA has only 500 instances and has a high prediction error for most design choices. Even using LM loss as the intermediate feature, while the prediction error is low, the fitting error is actually high (2.35%) as shown in §C.3, indicating randomness. On the other hand, MMLU and HellaSwag have large sample sizes – 5x larger than the other tasks. Given their low sample variance and noise around the prediction target, we expect that their errors are most representative of the true difficulty of predicting downstream performance, so even though taskCE and task loss perform comparably for several tasks, task loss is better considering task size. If the goal is to select a single design choice that works for all tasks, the design choice should be validated on tasks with a large number of instances.
Robustness of the ladder models.
In §D, we predict the task accuracies for a larger OLMo 2 target model: 32B-6T, using the same ladder models (0.45% of the target model compute). While the absolute errors are higher, they still follow the same trend as the smaller target models (lower variance tasks exhibit relatively lower prediction errors), indicating that the same ladder can reliably be used to estimate the performance of much larger models.
@@ 7 Related Work
Kaplan et al. (2020) and Hoffmann et al. (2022) were among the first to postulate the functional form of LM losses as a power function of model parameters and data size. The Chinchilla equation, in particular, has become the basis of many subsequent scaling law works (Muennighoff et al., 2023; Gadre et al., 2024; Zhang et al., 2024). This line of work focuses on predicting the LM loss (rather than downstream performance) on a held-out set with similar distribution as the pretraining data, but not on downstream tasks which is more important and challenging.
@@
Scaling laws for downstream tasks have been explored in Gadre et al. (2024). Instead of directly predicting accuracy for individual tasks, they compute the average top-1 error over 17 LLM-foundry (MosaicML, 2024) evaluation tasks as a function of cross entropy loss on C4 (Raffel et al., 2019). Dubey et al. (2024) uses a two-step prediction to first map the the training compute to the negative log-likelihood of the correct answer for a single task in an evaluation benchmark, and then relate the log-likelihood to the task accuracy. Unlike our work, which uses a fixed ladder of small models, they rely on first finding compute-optimal models. Chen et al. (2024) also employs a two-stage approach for predicting downstream performance, but uses the pre-training loss instead of a task-specific loss as the intermediate step. Isik et al. (2024) studies scaling laws in the finetuning setup, for machine translation tasks. Polo et al. (2025) leverages models of different families to predict downstream performance, but this leads to less accurate predictions for guiding pretraining development for a specific family of models.
@@ 8 Conclusion and Future Work
We develop model ladders and task scaling laws to predict downstream task performance of overtrained LMs (7B–4T and 13B–5T). Specifically, our contributions include predicting individual task performance for a range of tasks with multiple design choices; using a fixed set of ladder models (rather than using a large set of models to find compute-optimal models for each size); and using a small compute budget (1% of target compute). We also conduct variance analysis to determine task predictability, and provide recommendations for picking good design choices.
In future work, we hope to explore ways to improve task predictability, such as increasing the size of evaluation sets, alternative evaluation formats, reducing the impact of randomness, etc. We also hope to extend our prediction method to tasks in the multiple-choice (MC) format to more accurately reflect the capabilities of larger LMs. MC accuracy is harder to extrapolate from our smaller ladder models; we show some preliminary results in §B.2. Finally, we also hope to validate our method on larger models (70B parameters and beyond).